1.0 dm3 of an ideal gas at 100 kPa and 25 °C is heated to 50 °C at constant pressure. What is the new volume in dm3?




1
Expert's answer
2021-08-18T16:16:01-0400

"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"

"\frac{100000 \times 1}{273+25}=\frac{100000\times V_2}{273+50}"

"298V_2=323"

"V_2=323\/298=1.084dm^3"